OpenStudy (anonymous):

Can you think of a few examples of how we write mathematical expressions in real-life?

OpenStudy (anonymous):

im having trouble writing how i use them

OpenStudy (turingtest):

pretty much say anything you can think of relating two things by some numerical relation and it can be translated into algebra

OpenStudy (anonymous):

I am thinking of a number. If I add 2 to that number, I will get 5. What is the number? Although it may be fairly easy to guess that the number is 3, Let x be the number in my mind. Add 2 to x to get 5 Adding 2 to x to get 5 means that whatever x is, when I add 2 to x, it has to equal to 5 The equation is 2 + x = 5

OpenStudy (turingtest):

johnny is older than kate age of johnny=J age of kate=K translates to J>k

OpenStudy (anonymous):

Soon or later, all of us use the service of a taxi driver Taxi drivers usually charge a an initial fixed fee as part of using their services. Then, for each mileage, they charge a certain amount Say for instance, the initial fee is 4 dollars and each mileage cost 2 dollars The total cost can be modeled with an equation that is linear. Let y be the total cost Let N be number of mileage Total cost = 4 + cost for N miles Notice that cost for N miles = N×2 Therefore, y = 4 + N×2 Say for instance, a taxi driver takes you to a distance of 20 miles, how much money do you have to pay using y = 4 + N×2 ? When N = 20, Y = 4 + 20×2 = 4 + 40 = 44 dollars Now, let's ask the question the other way around! If you pay 60 dollars, how far did the taxi driver took you? This time y = 60 Replacing 60 into the equation gives you the following equation: 60 = 4 + N×2 It is not obvious to see that N = 28. That is why it is important to learn to solve linear equations!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome :)

OpenStudy (turingtest):

welcome :)

OpenStudy (turingtest):

I just went for a very simple example there are also classic problems like... I have a total of 20 coins they are all dimes and quarters that total $11.50 how many of each coin do I have? x=#dimes y=#quarters if you have 20 coins then x+y=20 and if the amount of money totals (money per coin)x(# that coin) then 0.1x+0.25y=11.50 you then have two equations x+y=20 0.1x+0.25y=11.50 you can use these two eqns to solve for the number of each coin, x an y